Grade 12Physical Chemistry

Calculate the mole fraction of ethyl alcohol and water in a solution containing 9.2 g of alcohol in 18 g of water

Gaveshk spolia
8 Years agoGrade 12

1 Answer

Gaveshk spolia
8 Years ago
Moles of ethanol =weight of ethanol /molar mass of ethanol =9.2/46=0.2Moles of water =weight of water/ molar mass of water=18/18=1Mole fraction of ethanol=0.2/0.2+1 =0.1661Mille fraction of water =1/1.2=0.8333